1. Assume that the demand curve is given by Q = 1000 – 0.25P. What is the inverse demand curve?

B) Using the inverse demand curve you solved for in 1, solve for the total revenue for this

Monopolist.

C) Using the total revenue curve you solved for in 2, solve for the marginal revenue curve.

D) Assume that Marginal costs are given by 100 + 2Q. What is the profit-maximizing quantity and

price for the monopolist?

E) Now, turn your attention to the Excel Spreadsheet. What is the Profit Maximizing quantity and

price for the monopolist?
Answer #1

Assume that the demand curve is given by Q = 1000 – 0.25P.

A) Inverse demand curve is given by

Q = 1000 - 0.25P

0.25P = 1000 - Q

P = 1000/0.25 - Q/0.25

P = 4000 - 4Q

B) Total revenue for this Monopolist is R = PQ

R = (4000 - 4Q)Q

R = 4000Q - 4Q^2

C) Marginal revenue curve is derivative of R

MR = dR/dQ

MR = 4000 - 8Q

D) Assume that Marginal costs are given by 100 + 2Q.

At profit maximizing quantity we have MR = MC

4000 - 8Q = 100 + 2Q

3900 = 10Q

Q = 390 units

P = 880 per unit

E) It is same as D)

Q = 390 units

P = 880 per unit

Both MR = MC = $880

Q MR MC
0 4000 100
100 3200 300
200 2400 500
300 1600 700
350 1200 800
390 880 880
400 800 900